OpenStudy (anonymous):

I need help finding the volume of a solid by rotating, one graph is y=0.5x^2 and the other is y=x. The region bounded by the two is rotated around the line x=3, i'm confused as to how to do this, any help would be nice.

OpenStudy (amistre64):

what parts are not confusing to you ... lets start there

OpenStudy (amistre64):

you thinking shells or washers?

OpenStudy (anonymous):

I was thinking washer, correct?

OpenStudy (amistre64):

either way, whatevers simplest for you.

OpenStudy (amistre64):

what are our points of intersection? 0,0 and what else?

OpenStudy (anonymous):

0,0 and 2,2

OpenStudy (amistre64):

so if we do washers, whats our formula for area of a washer?

OpenStudy (amistre64):



OpenStudy (anonymous):

\[\int\limits_{0}^{2} \Pi [(f(x)^2 - (g(x)^2]\]

OpenStudy (amistre64):

well, R^2 and r^2 they may not be singular functions

OpenStudy (amistre64):

 R=3-y r=3-sqrt(2y)